OpenStudy (anonymous):

find the volume of a solid with cross-sectional area (1/4x^2+4x+5) lying between the planes x=0 and x=0.5

OpenStudy (anonymous):

Integrate between 0 and 0.5

OpenStudy (anonymous):

shouldn't it be a double integral as it is a volume not an area

OpenStudy (anonymous):

Since this is adding up area, we'd only need single integral.The function represent the area

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

if yu were eistein+newton+gallileo..shouldnt yu know this??

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

finding the integral is a bit hard I guess. 4x^2+4x+5=4(x^2+x+5/4)= =4(1 + (x+1/2)^2)

OpenStudy (anonymous):

now you can see that the it looks nearly like 1/(1+x^2)

OpenStudy (anonymous):

do you see the solution now or should I still help?

OpenStudy (anonymous):

The person who asked the question is no longer around

OpenStudy (anonymous):

ok than I just worked with it for fun :-)

OpenStudy (anonymous):

wait i should try it out il then il tell you what i get...